OpenStudy (anonymous):

There is an inverse relationship between x and y. If y is sixteen when x is 4, what is x when y is 8? i want yall to tell me if this is right i think that the X is going to be 2

OpenStudy (amistre64):

y=k/x for some constant k 16 = k/4, k = 64 so our relationship is simply: y = 64/x, or x = 64/y in this case